# Neutrino Heating in 1D, 2D, and 3D core-collapse supernovae: characterizing the explosion of high-compactness stars

### v1

@@ Abstract
keyword1 – keyword2 – keyword3
@@ 1 Introduction
The explosion of core-collapse supernovae (CCSNe) is a long-standing problem in astrophysics. The very first simulations of CCSNe were carried out assuming spherical symmetry, simple prescriptions for the equation of state at high densities, and an approximate treatment of neutrino transport and interactions. Despite their simplicity, these pioneering models were the first to establish the role of neutrinos in driving the explosion (Colgate & White, 1966; Arnett, 1966; Bethe & Wilson, 1985). The tremendous advancements in computational resources in the last few decades allowed for more and more sophisticated simulations to become feasible (Herant et al., 1994; Miller et al., 1993; Fryer & Warren, 2002; Summa et al., 2018; Bruenn et al., 2020; Burrows et al., 2020; Nakamura et al., 2024; Kuroda et al., 2016). Moreover, the significant improvement in models for the nuclear equation of state (Lattimer & Swesty, 1991; Hempel & Schaffner-Bielich, 2010; Raduta et al., 2021; Fischer et al., 2017) and neutrino interactions (Bruenn, 1985; Hannestad & Raffelt, 1998; Horowitz, 2002; Buras et al., 2003; Burrows et al., 2006; Bollig et al., 2017; Horowitz et al., 2017) significantly reduced the uncertainties that enter into the simulations. However, several uncertainties are still affecting the models and their predictions. Therefore, the explosion of core-collapse supernovae remains a very complex phenomenon with many questions that need to be answered.
@@
$\xi_{M}=\dfrac{M/M_{\odot}}{R(M_{\rm bary}=M)/1000\,{\rm km}},$ (1)
where $R$ is the radius that encloses a mass $M_{\rm bary}=M$.
@@
Until recently (Janka, 2012), it was believed that all stars above $18-20M_{\odot}$ (i.e. with large iron cores and high compactness) would not explode, whereas all stars below that would yield a successful supernova. This was in part motivated by the fact that no type II-P SNe were observed for stars above $\gtrsim 20M_{\odot}$ (Smartt et al., 2009; Smartt, 2015), the so-called Red Supergiant Problem. However, this problem is now considered to be much less significant and can mostly be attributed to the large uncertainties in the imaging of the pre-SN progenitors (Beasor et al., 2024; Davies & Beasor, 2018, 2020). At the same time, current models have shown that the explodability has a much more complicated dependency on initial mass of the progenitor (Pejcha & Thompson, 2015; Sukhbold et al., 2016; Ebinger et al., 2019; Couch et al., 2020; Müller et al., 2016a; Boccioli et al., 2023; Fryer et al., 2022). Still, the fate of high-compactness stars is not well understood, and in some studies (Pejcha & Thompson, 2015; Sukhbold et al., 2016; Müller et al., 2016a; Ebinger et al., 2019; Fryer et al., 2022), these stars are predicted (or oftentimes assumed) to yield failed supernovae. In others (Couch et al., 2020; Boccioli et al., 2023; Wang et al., 2022), they are predicted to yield successful explosions instead. Given this tension, part of the motivation behind this study is to analyze the differences among these 1D studies, specifically with respect to the explosion of high-compactness progenitors.
All of the explodability studies mentioned above that, by definition, require hundreds of simulations, have been carried out using spherically symmetric (1D) models. The only exceptions are a recent study where 100 2D simulations were used (Vartanyan & Burrows, 2023; Wang et al., 2022; Tsang et al., 2022) and a less recent but more parametrized one (Nakamura et al., 2015), also in 2D.
@@
All of the 1D models used in the explodability studies mentioned above are characterized by different degrees of approximations. Some rely on simple semi-analytic models and others rely on self-consistent simulations where extra heating has been added. Ultimately, the explosion in all of these models is triggered by neutrinos, as also shown by all modern high-fidelity simulations (but see Soker (2010) for a semi-analytical model where neutrino heating is not the main cause of the explosion). Therefore, to study the explodability of high-compactness progenitors it is crucial to understand how neutrino heating changes with compactness.
@@
In the more recent past, there have been several multi-dimensional simulations that showed how high-compactness progenitors can produce successful explosions, often accompanied by the formation of large black holes around $\sim 20-30M_{\odot}$ (Müller et al., 2018; Powell et al., 2021; Sykes & Müller, 2024; Eggenberger Andersen et al., 2024; Chan et al., 2018; Burrows et al., 2023, 2024a). This however goes against what has often been assumed by previous studies (Fryer et al., 2012; Sukhbold et al., 2016; Ebinger et al., 2019; Pejcha & Thompson, 2015; O’Connor & Ott, 2011). This is particularly relevant for population synthesis (Iorio et al., 2023; Riley et al., 2022; Breivik et al., 2020; Eldridge et al., 2017; Belczynski et al., 2008; Kamlah et al., 2022; Mennekens & Vanbeveren, 2014; Izzard et al., 2004; Andrews et al., 2024) and galactic chemical evolution (GCE) simulations, (Diehl & Prantzos, 2023; Côté et al., 2017; Johnson & Weinberg, 2020; Andrews et al., 2017; Gjergo et al., 2023; Rybizki et al., 2017; Chen et al., 2023; Kubryk et al., 2015; Spitoni et al., 2020) both of which adopt prescriptions for the explodability of massive stars where high-compactness progenitors do not explode. Since the overwhelming majority of GCE and population synthesis calculations assume that high-compactness progenitors do not explode (but see Jost et al. (2024) for an exception), it is not clear how sensitive these models are to this assumption. Therefore, more work is needed to understand the fate of high-compactness stars.
@@
Given the rarity of these events, observational signatures of explosions of high-compactness stars are quite hard to predict. Since high-compactness progenitors tend to be high-mass stars, they are significantly disfavored by the initial mass function (IMF). Additionally, the stars with the largest compactness tend to form at metallicities much lower than solar and are therefore more abundant in distant galaxies, where faint supernovae are harder to detect. If the fallback onto the central black hole is significant, a high-compactness star might not produce any explosion at all (Summa et al., 2018; Kuroda et al., 2018; Walk et al., 2020), a weak explosion one or two orders of magnitude dimmer than regular type-II CCSNe (Eggenberger Andersen et al., 2024; Chan et al., 2018; Sykes & Müller, 2024), or a quite strong explosion instead (Burrows et al., 2024a; Sykes & Müller, 2024; Burrows et al., 2023). Given the few detailed studies on this subject and the differences among multi-dimensional simulations from different groups, this is still an open problem that requires deeper investigation. Another reason why predicting their observational signatures is challenging is that the uncertainties in the late phases of the explosion are quite large since long-term simulations are computationally very expensive (Vartanyan et al., 2024).
@@
In the next two years the ground-based Vera C. Rubin Observatory (Ivezić et al., 2019) will begin on-sky operations. Similarly, in the next two decades next-generation gravitational wave observatories Cosmic Explorer (CE) (Reitze et al., 2019; Evans et al., 2021) and Einstein Telescope (ET) (Punturo et al., 2010; Maggiore et al., 2020) will go online. This will provide an unprecedented number of observed light curves and detections of compact remnants, in particular black holes. Therefore, rarer events such as explosions of high-mass, low-metallicity progenitors, will become more likely to be detected, and better models are needed to understand their properties. Moreover, the much more detailed measurements of the black-hole mass distribution by CE and ET could be used to put constraints on which stars should explode and which ones should collapse into black holes. Specifically relevant for this work, it will become possible to constrain how rare are events where high-compactness stars explode and at the same time leave behind a black hole of $20-30M_{\odot}$.
@@
In this paper, we study how neutrino heating during the explosion of high-compactness progenitors from three different codes: GR1D, FLASH, and Fornax. We will analyze 1D, 1D+, 2D, and 3D simulations. For the remainder of this paper, we will refer to simulations in spherical symmetry where neutrino-driven convection is included using a parametric model as 1D+ simulations. In Section 2 we describe neutrino heating and neutrino-driven convection in the gain region, and describe how convection is modeled in 1D+ simulations. In Section 3 we compare 1D+, 2D and 3D models from different codes. In Section 4 we analyze the interplay between neutrino-driven convection and neutrino heating in 1D+ models as a function of compactness. In Section 5 we compare 1D+ models with other widely-used 1D models in the literature. Finally, in Section 6 we provide a summary and draw the main conclusions of this work.
@@ 2 Neutrinos in the gain region
To understand how the explosion of high-compactness progenitors develops, it is instructive to highlight how neutrinos directly (i.e. by being absorbed by the material close to the shock) and indirectly (by triggering convection) affect the revival of the shock. After the initial shock expansion, the shock stalls during the accretion phase due to a delicate balance between the energy lost by photodissociating the infalling material, and the energy gained due to neutrino absorption. The region where the energy gained by neutrino absorption is larger than the energy lost due to neutrino emission is called the gain region. During this stalled-shock phase, neutrino-driven convection plays a crucial role in further supporting the shock and increasing the time that matter spends in the gain region.
@@ 2.1 Neutrino-driven convection in the gain region
Neutrino-driven turbulent convection contributes to the overall energy budget in the gain region in two ways: (i) dissipation of turbulent kinetic energy into heat and (ii) transport of hot bubbles of material closer to the shock. These two tightly related effects can be quantitatively understood by analyzing the Reynolds decomposed Euler equations, as shown by Murphy & Meakin (2011); Murphy et al. (2013); Mabanta & Murphy (2018); Couch et al. (2020).
Figure 1: Left panels show the rate of generation of turbulent kinetic energy in the gain region, i.e. the integrand in Eq. (9), for 9 1D and 1D+ simulations of the progenitors listed in Table 1. The middle panels show the net neutrino heating in the gain region, i.e. the integrand in Eq. 8 for the same 1D+ models. The right panels show the net neutrino heating in the gain region for 1D simulations of the same progenitors used in the 1D+ simulations. Each row corresponds to a different time post bounce (tpb). Notice that the x-axis is a normalized radial coordinate within the gain region. One can see that the radial dependence of the specific neutrino heating is the same in 1D and 1D+ simulations. However, the size of the gain region (i.e. $r_{\rm shock}-r_{\rm gain}$), which can be inferred from the two leftmost panels in Figure 6, is larger in 1D+ simulations, and therefore neutrino heating (i.e Eq. 8) is larger in 1D+ simulations. As seen in the left panel, in 1D+ simulations the turbulent energy contribution is a bit smaller than the neutrino contribution, and very close to the shock the rate of generation of turbulent energy decreases faster than the neutrino heating.
With an appropriate closure relation, the Reynolds decomposed equations can also be used to include the effect of convection into spherically symmetric simulations, also referred to as 1D+ models. The closure is used to determine the rate of turbulence dissipation $\epsilon_{k}$, the Reynolds stress, and the internal energy flux. In this paper, we consider the STIR model, originally developed by Couch et al. (2020) and generalized to GR hydrodynamics by Boccioli et al. (2021), although similar models have also been developed (Mabanta & Murphy, 2018; Sasaki & Takiwaki, 2024). The closure utilized by Couch et al. (2020) is based on mixing length theory, i.e. the internal energy flux is treated as a diffusion term, the Reynolds stress is expressed in terms of the convective velocity, and the turbulence dissipation is expressed in terms of the mixing length. To properly account for all of the extra terms due to turbulence, the Euler equations should be decomposed in a background and perturbed component, and more details on this can be found in the original papers (Murphy & Meakin, 2011; Murphy et al., 2013; Mabanta & Murphy, 2018; Couch et al., 2020). Here, we simply show the evolution equation for the turbulent energy in the STIR model:
@@
$F_{K}=-\alpha_{\rm MLT}D_{K}\rho\nabla v_{\rm turb}^{2}=-\alpha_{\rm MLT}\alpha_{\rm D}v_{\rm turb}\Lambda_{\rm MLT}\rho\nabla v_{\rm turb}^{2},$ (6)
where $v_{\rm turb}$ is the convective velocity, $D_{K}$ is a diffusion coefficient, $\Lambda_{\rm MLT}$ is the mixing length, and $\alpha_{\rm MLT}$ is a parameter of $\mathcal{O}(1)$ that effectively controls the strength of convection. The other parameter $\alpha_{\rm D}$ is instead kept constant at a value of $1/6$ (see Müller et al. (2016b); Couch et al. (2020)). The terms $\rho\epsilon_{g}$ and $\rho\epsilon_{k}$ are the rates of generation and dissipation of turbulent kinetic energy per volume, respectively. The turbulent kinetic energy can be generated via shear or buoyancy (i.e. $W_{s}$ and $W_{b}$ terms in Eq. (4)). In the full set of hydrodynamic equations, the term $\epsilon_{k}$ is added as a positive source term to the RHS of the evolution equation for the internal energy. However, as pointed out by Gogilashvili et al. (2023a), it is the term $W_{b}$ (i.e. the term responsible for generating the convection) that enters the equation for the total energy.
It should be highlighted that in the literature the shear term is usually neglected since it is quite small, but is included in STIR. Therefore, it is the term $\rho\epsilon_{g}=W_{s}+W_{b}$ that is responsible for generating turbulence. This can be more clearly seen if one considers the equation for the total energy, which includes the contribution from internal, kinetic, and turbulent energy:
@@
where $F_{e}$ is a term analogous to $F_{K}$, where instead of the gradient of $v_{\rm turb}^{2}$ one has the gradient of internal energy. As mentioned above, the main effects of convection are dissipation of turbulent kinetic energy into heat (i.e. $\rho\epsilon_{g}$) and transport of energy within the gain region (i.e. $\nabla\cdot(F_{e}+F_{K})$). In addition to the turbulent pressure $P_{\rm turb}$, which is quite small, these are the two most important extra terms added to the spherically symmetric Euler equations.

### v2

@@ Abstract
transients: supernovae – stars: massive – neutrinos
@@ 1 Introduction
The explosion of core-collapse supernovae (CCSNe) is a long-standing problem in astrophysics. The very first simulations of CCSNe were carried out assuming spherical symmetry, simple prescriptions for the equation of state at high densities, and an approximate treatment of neutrino transport and interactions. Despite their simplicity, these pioneering models were the first to establish the role of neutrinos in driving the explosion (Colgate & White, 1966; Arnett, 1966; Bethe & Wilson, 1985). The tremendous advancements in computational resources in the last few decades allowed for more and more sophisticated simulations to become feasible (Herant et al., 1994; Miller et al., 1993; Fryer & Warren, 2002; Kuroda et al., 2016; Summa et al., 2018; Bruenn et al., 2020; Burrows et al., 2020; Nakamura et al., 2025). Moreover, the significant improvement in models for the nuclear equation of state (Lattimer & Swesty, 1991; Hempel & Schaffner-Bielich, 2010; Fischer et al., 2017; Raduta et al., 2021) and neutrino interactions (Bruenn, 1985; Hannestad & Raffelt, 1998; Horowitz, 2002; Buras et al., 2003; Burrows et al., 2006; Bollig et al., 2017; Horowitz et al., 2017) significantly reduced the uncertainties that enter into the simulations. However, several uncertainties are still affecting the models and their predictions. Therefore, the explosion of core-collapse supernovae remains a very complex phenomenon with many questions that need to be answered.
@@
$\xi_{M}=\dfrac{M/M_{\odot}}{R(M)/1000\,{\rm km}},$ (1)
where $R(M)$ is the radius that encloses a baryonic mass $M$, with $M$ typically being chosen to be between 1.75 $M_{\odot}$ and 2.5 $M_{\odot}$. In the vast majority of this paper, we will be using $M=2.0\,M_{\odot}$.
@@
Until recently (Janka, 2012), it was believed that all stars above $18-20M_{\odot}$ (i.e. with large iron cores and high compactness) would not explode, whereas all stars below that would yield a successful supernova. This was in part motivated by the fact that no type II-P SNe were observed for stars above $\gtrsim 20M_{\odot}$ (Smartt et al., 2009; Smartt, 2015), the so-called Red Supergiant Problem. However, this problem is now considered to be much less significant and can mostly be attributed to the large uncertainties in the imaging of the pre-SN progenitors (Davies & Beasor, 2018, 2020; Beasor et al., 2025). At the same time, current models have shown that the explodability has a much more complicated dependency on the initial mass of the progenitor (Pejcha & Thompson, 2015; Sukhbold et al., 2016; Müller et al., 2016a; Ebinger et al., 2019; Couch et al., 2020; Boccioli et al., 2023; Fryer et al., 2022). Still, the fate of high-compactness stars is not well understood, and in some studies (Pejcha & Thompson, 2015; Sukhbold et al., 2016; Müller et al., 2016a; Ebinger et al., 2019; Fryer et al., 2022), these stars are predicted (or oftentimes assumed) to yield failed supernovae. In others (Couch et al., 2020; Boccioli et al., 2023; Wang et al., 2022), they are predicted to yield successful explosions instead. Given this tension, part of the motivation behind this study is to analyze the differences among these studies, specifically with respect to the explosion of high-compactness progenitors.
All of the explodability studies mentioned above that, by definition, require hundreds of simulations, have been carried out using spherically symmetric (1D) models. The only exceptions are two sets of hundreds of 2D simulations by Nakamura et al. (2015) and Vartanyan & Burrows (2023). The former study involved $\sim 350$ simulations, where, however, the transport of electron-type neutrinos was described using the simple (but computationally efficient) Isotropic Diffusion Source Approximation (Liebendörfer et al., 2009), whereas heavy lepton neutrinos were treated using an even more simplified approach following a leakage scheme. Moreover, the ray-by-ray approximation was adopted, i.e. neglecting non-radial effects in the neutrino transport. The latter study is more recent, was carried out for 100 simulations, and uses a state-of-the-art M1 transport scheme to solve for the transport of neutrinos. The explodability of those 100 2D simulations was then analyzed by Wang et al. (2022) and Tsang et al. (2022).
@@
All of the 1D models used in the explodability studies mentioned above are characterized by different degrees of approximations. Some rely on simple semi-analytic models and others rely on self-consistent simulations where extra heating has been added. Ultimately, the explosion in all of these models is triggered by neutrinos, as also shown by all modern high-fidelity simulations. Therefore, to study the explodability of high-compactness progenitors, it is crucial to understand how neutrino heating changes with compactness.
@@
In the more recent past, there have been several multi-dimensional simulations that showed how high-compactness progenitors can produce successful explosions, often accompanied by the formation of large black holes around $\sim 20-30M_{\odot}$ (Müller et al., 2018; Chan et al., 2018; Powell et al., 2021; Sykes & Müller, 2025; Burrows et al., 2023; Eggenberger Andersen et al., 2025; Burrows et al., 2024a). This however goes against what has often been assumed by previous studies (O’Connor & Ott, 2011; Fryer et al., 2012; Pejcha & Thompson, 2015; Sukhbold et al., 2016; Ebinger et al., 2019). This is particularly relevant for population synthesis (Izzard et al., 2004; Belczynski et al., 2008; Mennekens & Vanbeveren, 2014; Eldridge et al., 2017; Breivik et al., 2020; Riley et al., 2022; Kamlah et al., 2022; Iorio et al., 2023; Andrews et al., 2024) and galactic chemical evolution (GCE) simulations, (Kubryk et al., 2015; Côté et al., 2017; Andrews et al., 2017; Rybizki et al., 2017; Spitoni et al., 2020; Johnson & Weinberg, 2020; Gjergo et al., 2023; Chen et al., 2023; Diehl & Prantzos, 2020) both of which adopt prescriptions for the explodability of massive stars where high-compactness progenitors do not explode. Since the overwhelming majority of GCE and population synthesis calculations assume that high-compactness progenitors do not explode (but see Jost et al. (2025) for an exception), it is not clear how sensitive these models are to this assumption. Therefore, more work is needed to understand the fate of high-compactness stars.
@@
Given the rarity of these events, observational signatures of explosions of high-compactness stars are quite hard to predict. Since high-compactness progenitors tend to be high-mass stars, they are significantly disfavored by the initial mass function (IMF). Additionally, the stars with the largest compactness tend to form at metallicities much lower than solar and are therefore more abundant in distant galaxies, where faint supernovae are harder to detect. If the fallback onto the central black hole is significant, a high-compactness star might produce no explosion at all (Summa et al., 2018; Kuroda et al., 2018; Walk et al., 2020), a weak explosion one or two orders of magnitude dimmer than regular type-II CCSNe (Chan et al., 2018; Eggenberger Andersen et al., 2025; Sykes & Müller, 2025), or a quite strong explosion instead (Burrows et al., 2023, 2024a; Sykes & Müller, 2025). Given the few detailed studies on this subject and the differences among multi-dimensional simulations from different groups, this is still an open problem that requires deeper investigation. Another reason why predicting their observational signatures is challenging is that the uncertainties in the late phases of the explosion are quite large since long-term simulations are computationally very expensive (Vartanyan et al., 2025).
@@
In the next two years the ground-based Vera C. Rubin Observatory (Ivezić et al., 2019) will begin on-sky operations. Similarly, in the next two decades, next-generation gravitational wave observatories such as Cosmic Explorer (CE) (Reitze et al., 2019; Evans et al., 2021) and Einstein Telescope (ET) (Punturo et al., 2010; Maggiore et al., 2020) will go online. This will provide an unprecedented number of observed light curves and detections of compact remnants, in particular black holes. Therefore, rarer events such as explosions of high-mass, low-metallicity progenitors, will become more likely to be detected, and better models are needed to understand their properties. Moreover, the much more detailed measurements of the black-hole mass distribution by CE and ET could be used to put constraints on which stars should explode and which ones should collapse into black holes. Specifically relevant for this work, it will become possible to constrain how rare are events where high-compactness stars explode and at the same time leave behind a black hole of $20-30M_{\odot}$.
@@
In this paper, we study how neutrino heating develops during the explosion (with a particular focus on high-compactness progenitors) in three different codes: GR1D, FLASH, and Fornax. We will analyze 1D, 1D+, 2D, and 3D simulations. For the remainder of this paper, we will refer to simulations in spherical symmetry where neutrino-driven convection is included using a parametric model as 1D+ simulations. In Section 2, we describe neutrino heating and neutrino-driven convection in the gain region and describe how convection is modeled in 1D+ simulations. In Section 3 we compare 1D+, 2D and 3D models from different codes. In Section 4 we analyze the interplay between neutrino-driven convection and neutrino heating in 1D+ models as a function of compactness. In Section 5 we compare 1D+ models with other widely-used 1D models in the literature. Finally, in Section 6 we provide a summary and draw the main conclusions of this work.
@@ 2 Neutrinos in the gain region
To understand how the explosion of high-compactness progenitors develops, it is instructive to highlight how neutrinos directly (i.e. by being absorbed by the material close to the shock) and indirectly (by triggering convection) affect the revival of the shock. After the initial shock expansion, the shock stalls during the accretion phase due to a delicate balance between the energy lost by photodissociating the infalling material, and the energy gained due to neutrino absorption. The region where the energy gained by neutrino absorption is larger than the energy lost due to neutrino emission is called the gain region. During this stalled-shock phase, neutrino-driven convection plays a crucial role in further supporting the shock and increasing the time that matter spends in the gain region. In practice, throughout this paper, we define the gain region to be the region where net neutrino heating is positive, density is below $3\times 10^{10}\ {\rm g/cm^{3}}$, and specific entropy per baryon is above $6$. We will come back to the importance of these cuts in density and entropy in Section 3.
@@ 2.1 Neutrino-driven convection in the gain region
Neutrino-driven turbulent convection contributes to the overall energy budget in the gain region in two ways: (i) dissipation of turbulent kinetic energy into heat and (ii) transport of hot bubbles of material closer to the shock. These two tightly related effects can be quantitatively understood by analyzing the Reynolds-decomposed Euler equations, which separate each variable into a mean background component and a perturbed (turbulent) one (Murphy & Meakin, 2011; Murphy et al., 2013; Mabanta & Murphy, 2018; Couch et al., 2020).
Figure 1: The first four rows of this Figure show the 9 1D+ simulations for the progenitors listed in Table 1 color-coded by compactness. The last row instead show the same 9 progenitors but refers to a simple 1D simulation. Each column shows snapshots at three different times post-bounce. Notice that all quantities have been multiplied by the volume to highlight the contribution to the total net heating at each radius. The x-axis shows the dimensionless radial coordinate within the gain region. This choice is motivated by the need to show different simulations with different shock radii and radial extents of the gain region in the same canvas. The total net neutrino heating rate and rate of turbulent energy generation integrated over the gain region for these same simulations is shown in Section 4. Notice the different ranges for the y-axis. Only the fourth and fifth rows have the same.
With an appropriate closure relation, the Reynolds decomposed equations can also be used to include the effect of convection into spherically symmetric simulations, also referred to as 1D+ models. The closure is used to determine the rate of turbulent energy dissipation $\epsilon_{k}$, the Reynolds stress, and the internal energy flux. In this paper, we consider the STIR model, originally developed by Couch et al. (2020) and generalized to GR hydrodynamics by Boccioli et al. (2021), although similar models have also been developed (Mabanta & Murphy, 2018; Sasaki & Takiwaki, 2024). The closure utilized by Couch et al. (2020) is based on mixing length theory, i.e. the internal energy flux is treated as a diffusion term, the Reynolds stress is expressed in terms of the convective velocity, and the turbulent energy dissipation is expressed in terms of the mixing length. To properly account for all of the extra terms due to turbulence, the Euler equations should be decomposed in a background and perturbed component, and more details on this can be found in the original papers (Murphy & Meakin, 2011; Murphy et al., 2013; Mabanta & Murphy, 2018; Couch et al., 2020). Here, we simply show the evolution equation for the turbulent energy in the STIR model:
@@
$F_{K}=-\alpha_{\rm MLT}D_{K}\rho\nabla v_{\rm turb}^{2}=-\alpha_{{\rm D}_{K}}v_{\rm turb}\Lambda_{\rm MLT}\rho\nabla v_{\rm turb}^{2}.$ (6)
In the above equations, $r$ is the radius, $\rho$ is the density, $v_{\rm turb}$ is the turbulent (or convective) velocity, $v_{r}$ is the radial velocity, $\Lambda_{\rm MLT}$ is the mixing length, $\omega_{\rm BV}$ is the Brunt-Väisälä/ frequency, $P$ is the pressure. We also define the quantities $\rho\epsilon_{g}$ and $\rho\epsilon_{k}$ to be the rates of generation and dissipation of turbulent kinetic energy per volume, respectively. The term $F_{K}$ indicates the turbulent energy flux, and in this model, it is treated as a diffusion term, where $D_{K}$ is its respective diffusion coefficient.
The most important parameter in this model is $\alpha_{\rm MLT}$, a parameter of $\mathcal{O}(1)$ that effectively controls the strength of convection. The other parameter that appears in the above equations is $\alpha_{{\rm D}_{K}}$, which controls how fast the diffusion of turbulent energy is. Similar terms are present in the equations for internal energy, electron fraction, and neutrino energy density, each with their own free parameter analogous to $\alpha_{{\rm D}_{K}}$. In accordance with several other studies in the literature (Müller et al., 2016b; Couch et al., 2020), all of these parameters are kept constant at a value of $1/6$. However, see Sasaki & Takiwaki (2024) for a study where the strength of diffusion is changed independently for different quantities.
As seen in Eq. (4), the turbulent kinetic energy can be generated via shear ($W_{s}$) or buoyancy ($W_{b}$). In the full set of hydrodynamic equations, the term $\epsilon_{k}$ is added as a positive source term to the RHS of the evolution equation for the internal energy. However, as pointed out by Gogilashvili et al. (2024), it is the term $W_{b}$ (i.e. the term responsible for generating the convection) that enters the equation for the total energy.
It should be highlighted that in the literature, the shear term is usually neglected since it is quite small, but is included in STIR. Therefore, it is the term $\rho\epsilon_{g}=W_{s}+W_{b}$ that is responsible for generating turbulent energy. This can be more clearly seen if one considers the equation for the total energy, which includes the contribution from internal, kinetic, and turbulent energy:
@@
where $F_{e}$ is a term analogous to $F_{K}$, where instead of the gradient of $v_{\rm turb}^{2}$ one has the gradient of internal energy, $g$ the local gravitational acceleration, and $\dot{q}_{\nu}=\rho\dot{\epsilon}_{\nu}$ is the net neutrino heating rate. As mentioned above, the main effects of convection are the generation and dissipation of turbulent kinetic energy into heat (i.e. $\rho\epsilon_{g}$ and $\rho\epsilon_{k}$) and transport of energy within the gain region (i.e. $\nabla\cdot(F_{e}+F_{K})$). In addition to the turbulent pressure $P_{\rm turb}$, which is quite small, these are the most important extra terms added to the spherically symmetric Euler equations.